Write in trig form: -6 + 6sqrt3i
Answer by stanbon(75887)

Write in trig form: -6 + 6sqrt3i
-----------------------------
Conversion formulas:
r = sqrt(a^2+b^2)
theta = tan^-1(b/a)
-------------------
Your Problem:
r = sqrt((-6)^2+(6sqrt3)^2)= sqrt(36+108)= 12
-----
theta = tan^-1(6sqrt3/-6) = tan^-1(-sqrt3)=(2/3)pi
--------------
Trig form:
-6+6sqrt3i = 12(cos((2/3)pi) + isin((2/3)pi)
============
Cheers,
Stan H.